# Cybersecurity Defenses: Exploration of CVE Types through Attack Descriptions

### v1

@@ II-C Designing VULDAT Architecture
Furthermore, multi-qa-distilbert-cos-v1 and all-distilroberta-v1, which are based on the DistilBERT and DistilRoBERTa architectures, respectively. Using the MPNet architecture, the multi-qa-mpnet-base-dot-v1 and all-mpnet-base-v2. Lastly, the MiniLM architecture, all-MiniLM-L6-v2, paraphrase-multilingual-MiniLM-L12-v2, and paraphrase-MiniLM-L6-v2 models, are designed for use in multilingual paraphrasing tasks. The sentence transformer models generate embeddings for the AT and the CVE descriptions. Then, we calculate a cosine similarity (Equation (II-D)) score between these vectors, which measures how closely the CVE descriptions match the AT text. Thus, the similarity score helps identify and assign the most relevant CVEs to each AT.
@@ II-D Compute Similarity Score
We apply our model to associate an AT with CVE report vectors by leveraging their cosine similarity scores. The output CVE vectors are ranked based on their similarity to the AT vector.
$Sim(\vec{p}\cdot\vec{q})=\frac{\vec{p}\cdot\vec{q}}{\|\vec{p}\|\cdot\|\vec{q}\|}=\frac{\sum_{i=1}^{n}p_{i}q_{i}}{\sqrt{\sum_{i=1}^{n}p_{i}^{2}}\cdot\sqrt{\sum_{i=1}^{n}q_{i}^{2}}}$
The top N CVE reports with a similarity score greater than 60% are then associated with the AT vector. This set of CVE reports, named the Detection List ($\mathcal{L}$), represents the most relevant vulnerabilities linked to the AT.
@@ III RESULTS
RQ2: How many CVE issues VULDAT correctly detect from an attack text?
To answer RQ2, we compute the Jaccard Similarity, Mapping Accuracy, and Detection Accuracy for each AT text. The Jaccard Similarity evaluates the similarity of two sets: the detected set ($\mathcal{L}_{a}$) by VULDAT and actual sets ($\mathcal{M}_{a}$). It measures the ratio of the intersection of these sets to their union, $\mathrm{Jaccard\,Similarity}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{|\mathcal{L}_{a}\cup\mathcal{M}_{a}|}$. Mapping Accuracy is the ratio of the intersection of the detected set ($\mathcal{L}_{a}$) and the actual set ($\mathcal{M}_{a}$) to the size of the actual set ($\mathcal{M}_{a}$), $\mathrm{Mapping\,Accuracy}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{{|\mathcal{M}_{a}|}}$. Detection Accuracy is the ratio of the intersection of the detected set ($\mathcal{L}_{a}$) and the actual set ($\mathcal{M}_{a}$) to the size of the detected set ($\mathcal{L}_{a}$), $\mathrm{Detection\,Accuracy}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{{|\mathcal{L}_{a}|}}$. A higher Detection Accuracy implies more accurate identifications by VULDAT.
@@ IV RELATED WORK
The Cve2att&ck  [21] model uses BERT-based [18] language models and classic models for machine learning to automatically link CVE reports with techniques that are able to link CVE issues with the MITRE ATT&CK Enterprise Matrix techniques by utilizing the textual description present in CVE metadata. The CVE Transformer (CVET)  [22] model uses RoBERTa [19] to link a CVE to ten tactics from ATT&CK Enterprise Matrix. Kuppa et al. [23] introduced the Multi-Head Joint Embedding Neural Network model to automatically map CVE issues to ATT&CK techniques. The suggested representation finds particular regular expressions in the current threat reports and then compares the ATT&CK technique vectors with the text description found in the CVE metadata by measuring their similarity using the cosine distance. SMET  [24] is an automatic mapping tool that matches CVE entries to ATT&CK techniques using textual similarity. It uses the ATT&CK BERT model, trained on the SIAMESE network  [25], to determine the semantic similarity between attack actions.

### v2

@@ II-C Designing VULDAT Architecture
Furthermore, multi-qa-distilbert-cos-v1 and all-distilroberta-v1, which are based on the DistilBERT and DistilRoBERTa architectures, respectively. Using the MPNet architecture, the multi-qa-mpnet-base-dot-v1 and all-mpnet-base-v2. Lastly, the MiniLM architecture, all-MiniLM-L6-v2, paraphrase-multilingual-MiniLM-L12-v2, and paraphrase-MiniLM-L6-v2 models, are designed for use in multilingual paraphrasing tasks. The sentence transformer models generate embeddings for the AT and the CVE descriptions. Then, we calculate a cosine similarity score between these vectors, which measures how closely the CVE descriptions match the AT text. Thus, the similarity score helps identify and assign the most relevant CVEs to each AT.
@@ II-D Compute Similarity Score
We apply our model to associate an AT with CVE report vectors by leveraging their cosine similarity scores. The output CVE vectors are ranked based on their similarity to the AT vector. The top N CVE reports with a similarity score greater than 60% are then associated with the AT vector. This set of CVE reports, named the Detection List ($\mathcal{L}$), represents the most relevant vulnerabilities linked to the AT.
@@ III Results
RQ2: How many CVE issues VULDAT correctly detect from an attack text? To answer RQ2, we compute the Jaccard Similarity, Mapping Accuracy, and Detection Accuracy for each AT text. The Jaccard Similarity evaluates the similarity of two sets: the detected set ($\mathcal{L}_{a}$) by VULDAT and actual sets ($\mathcal{M}_{a}$). It measures the ratio of the intersection of these sets to their union, $\mathrm{Jaccard\,Similarity}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{|\mathcal{L}_{a}\cup\mathcal{M}_{a}|}$. Mapping Accuracy is the ratio of the intersection of the detected set ($\mathcal{L}_{a}$) and the actual set ($\mathcal{M}_{a}$) to the size of the actual set ($\mathcal{M}_{a}$), $\mathrm{Mapping\,Accuracy}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{{|\mathcal{M}_{a}|}}$. Detection Accuracy is the ratio of the intersection of the detected set ($\mathcal{L}_{a}$) and the actual set ($\mathcal{M}_{a}$) to the size of the detected set ($\mathcal{L}_{a}$), $\mathrm{Detection\,Accuracy}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{{|\mathcal{L}_{a}|}}$. A higher Detection Accuracy implies more accurate identifications by VULDAT.
@@ IV Related Work
The Cve2att&ck  [21] model uses BERT-based [18] language models and classic models for machine learning to automatically link CVE reports with techniques that are able to link CVE issues with the MITRE ATT&CK Enterprise Matrix techniques by utilizing the textual description present in CVE metadata. The CVE Transformer (CVET)  [22] model uses RoBERTa [19] to link a CVE to ten tactics from ATT&CK Enterprise Matrix. Kuppa et al. [23] introduced the Multi-Head Joint Embedding Neural Network model to automatically map CVE issues to ATT&CK techniques. The suggested representation finds particular regular expressions in the current threat reports and then compares the ATT&CK technique vectors with the text description found in the CVE metadata by measuring their similarity using the cosine distance. SMET  [24] is an automatic mapping tool that matches CVE entries to ATT&CK techniques using textual similarity. It uses the ATT&CK BERT model, trained on the SIAMESE network, to determine the semantic similarity between attack actions.
@@ VI Conclusion
VII Acknowledgements
The first author thanks the CSLab at the Free University of Bozen-Bolzano for supporting this work under project no. EFRE1039 in the EFRE-FESR 2021-2027 program.